Hi
- If a question gives an exponetial or logarithmic equation and asks you to find the x and y-intercepts how would you find them?
For example:
y = 2^(x+1) - 4
y = log2(x+5)
Thankyou!

Answer by ikleyn(52847)

.
1.  For the equation  y = 2^(x+1) - 4

        y-interception  is the value of "y" at  x= 0,   which in this case is 2-4 = -2.

        x-interception  is the value of "x" where y = y(x) = 0.   

              In this case it is the root of the equation  2^(x+1) - 4 = 0, which is x = 1.